Algebra

The number of possible many to one function from A={6,36} to B={1,2,3,4,5} is

Ashlesha Rathi
12 Years agoGrade

1 Answer

surya
12 Years ago
no.of functions=5^2=25
no.of one one functions are 5P2=(5!)/(5-2)!=20
sp no.of  many one functions = 25-20=5